OpenStudy (anonymous):

Use the functions to answer the question. f(x) = 3x - 9; g(x) = x - 4 What is the value of (f o circle g)(5)? (Points : 4)

OpenStudy (anonymous):

f(x) = -x^2/3 + 2x + 7 x = 2/2(-1/3) = 2/(-2/3)= (2/1)(3/2)= 3 maximum (-x^2 + 6x + 21 /3) = ((-x^2 + 6x + 9) + 21 + 9)/3) = (-(x + 3)^2 + 30)/3 = -(x + 3)^2 + 10 y = 10

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

- 6 5 6 9

OpenStudy (anonymous):

(3x-2) squared is 9x^2 - 12x + 4

OpenStudy (anonymous):

did that help?

OpenStudy (anonymous):

no -_- lol, i am sorry

OpenStudy (shamil98):

(f o g) = you input the equation of g into f for example if f = x^2 and g = (x-3) you would put g into f like so (x-3)^2

OpenStudy (shamil98):

(f o circle g)(5) just means you input 5 into the equation of g, then substitute g into the equation for f.

OpenStudy (shamil98):

If g = x- 5 and f = 2x for g(5) g = 5-5 f = 2(0) f = 0 like so, do you understand?

OpenStudy (anonymous):

yes...